OpenStudy (anonymous):

write and equation of a line that includes the point ( - 4, 0 ) and has a slope of - 3 / 5 in slope - intercept form

OpenStudy (gabebae):

Do you know what slope intercept form is?

OpenStudy (anonymous):

a) y= -3/5x-12/5 b) y= -3/5x -12 c) y= -3/5x -5 d) y=-3/5x+12/5

OpenStudy (anonymous):

Yeah its y=mx+b

OpenStudy (gabebae):

do you know what the letters stand for?

OpenStudy (anonymous):

y is just y m is supposed to be the fraction/number that tells how the slope moves in a positive or negative slope x is just x and b us your starting point on the y axis

OpenStudy (gabebae):

I would go with A

OpenStudy (anonymous):

a is correct y= - 3 / 5x - 12/5

OpenStudy (gabebae):

good